I need help solving this system of equations by the substitution method:
y=2x-5
4x-y=7

Answer by askmemath(368)

y=2x-5 (1)
4x-y=7 (2)
From (1) we know that y = 2x-5
Put this back in (2)
4x-(2x-5) = 7
4x-2x+5 = 7
2x+5 = 7
Subtracting 5 on both sides
2x = 2
Dividing by 2 throughout
X = 1
which means y = 2(1)-5 = -3